A) For the reaction 2CH4(g)⇌C2H2(g)+3H2(g) Kc=0.150 mol2 L−2 at 1681 C. What is KP for the reaction at this temperature?

B) For the reaction N2(g)+3H2(g)⇌2NH3(g) KP=4.90×10−3 bar−2 at 289 C. What is Kc for the reaction at this temperature?
Answer #1

N 2C+y = Categy + 3Heady Kc = 0.150 molī? T= 1681+273 = 1954k An= product moles - Reactant molu = 4-2=2 R = 8.314 I Imol = 0